OpenStudy (anonymous):

Express 5^3 = x as a logarithmic equation. log53 = x log35 = x log3x = 5 log5x = 3 @satellite73

hartnn (hartnn):

\(\Large a^b=c \implies \log_ac=b\)

hartnn (hartnn):

or take log on both sides first

hartnn (hartnn):

then use \(\Large \dfrac{\log b}{\log a} = \log _ab\)